# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 4 Problem Classes & Instances
Figure 2: The figures show the number of variables against the density, i.e., the fraction of non-zero coefficients in the objective and constraint functions, of the various problem instances provided in the QOBLIB repository [193] for MIP (top) and QUBO (bottom) formulations of the problem. The legend entries for the various problem classes name the class itself and, in brackets, give the form of the (objective/constraint) function with L and Q denoting linear and quadratic, respectively. Furthermore, the size of the individual points is directly related to the coefficient range of the respective problem instance per problem class–see legend in the center right. The dashed lines illustrate the separation between sparse and dense instances as described in Table 2. Finally, the inset plots illustrate the instances in which the optimal solution is known. It should be noted that while the MIP plot shows all problem instances, the QUBO plot only shows those instances that we could transform into QUBO models with less than $1.6\cdot 10^{5}$ variables, cf. Table 2. The respective model translation is done for Birkhoff, Steiner, Sports, Network, Routing, and Topology with Qiskit Optimization 0.6.1 [192] using the default parameters. Marketsplit, LABS, Portfolio, and Independent Set can be trivially converted to QUBOs using ZIMPL 3.6.1 [134], since there are no slack variables to consider. Note that for Portfolio, this shows the density of the constraints matrix, the quadratic objective is always dense.
@@
The number of decision variables for which MIP or QUBO formulations of the selected instances become difficult ranges from less than 100 to 45,000, and is therefore within reach for testing algorithms on near-term quantum hardware. The constraint coefficients ranges lie, depending on the problem, between $1$ and $\approx 10^{7}$, assuming that the coefficient values are scaled to integers, as we do throughout the paper. This can pose another challenge for quantum computers as they need to be able to represent the problem data in the necessary resolution.
@@
Figure 2 provides a more detailed overview of the instances provided for the various problem classes complementing Table 2. For MIP and QUBO formulations of the considered problem instances, it illustrates the number of variables, the problem density, the coefficient range, and whether the optimal solution is known. Please note that if multiple MIP and QUBO models are available (see [193]), we choose to represent the one resulting in the smallest number of variables. We would further like to point out that the QUBO illustration only shows those instances that could be converted from an MIP model into a QUBO model with less than $1.6\cdot 10^{5}$ variables. As the plots show, the larger the number of variables, the more likely it is that we do not know the optimal solution. Finally, it can be seen that Market Split and LABS are the two problem classes for which difficult instances are found at the smallest sizes. Independent Set also has hard instances at relatively small sizes; these instances are also sparse and come with a small coefficient range. For these reasons, these three problem classes seem the most amenable for approaching with a quantum algorithm on near-term hardware.
@@
In the remainder of this section, we discuss each problem class, describe the specific instances included, and present baseline results computed with classical optimization solvers. Of course, all instances, used models, and solutions are also made accessible in the QOBLIB repository [193]. We do not claim that our baseline results are the best solutions one can find with a classical computer, but they should provide an indication of what is possible using general solvers without putting an extensive effort into fine-tuning. Moreover, the provided MIP and QUBO formulations were not subject to significant optimization efforts but are intended to serve as illustrations. Our goal is to provide a framework to track the progress towards quantum advantage in optimization and have an initial baseline to start with.
@@ 4.1.1 Background
The market split instances were explicitly designed to be hard to solve using classical methods. Since their introduction by [81] in 1999, the progress in solving them has been everything but spectacular. The time complexity for enumeration is $\mathcal{O}(2^{n/2}\cdot(n/4))$ [215]. Several other methods have been tried, including smart enumeration, branch-and-cut, and lattice point enumeration with basis reduction [5, 240, 235].
Why is it interesting? Market split problems are a type of a multi-dimensional subset-sum problem. This problem class contains various instances for which it is challenging to find feasible solutions with classical digital systems–already for small system sizes. More specifically, there exist instances with less than $100$ binary variables and $8$ constraints that cannot be solved. Applications of the market split problem could be interesting in various industries, such as the energy sector, e.g., to help derive optimal pricing strategies subject to customer acquisition, retention or competition constraints, and regional regulatory policies [89, 74].
@@ 4.1.2 Problem statement
where a feasible solution exists if the global minimum is zero. With careful choices of the coefficients, this problem becomes hard to solve, e.g., for numbers $m\in\mathbb{N}_{+}$, $D\in\mathbb{N}$, set $n=10(m-1)$, $I=\{1,\ldots m\},J=\{1,\ldots,n\}$ and $a_{ij}\in\{0,\ldots,D-1\},b_{i}=\left\lfloor\frac{1}{2}\sum_{j\in J}a_{ij}\right\rfloor$ [81].
The best results so far have been achieved using forms of lattice enumeration with basis reduction. [240] report solutions up to $m=9$. Commercial MIP solvers are typically more effective on the lattice reformulation introduced in [5], but often already struggle for $m\geq 7$. To the best of our knowledge, no solutions for instances with $m\geq 10$ have been reported in the literature so far.
@@ 4.1.3 Instances
For this benchmark, we generated four random instances each for $m=\{3,\ldots,15\}$, and $D=\{50,100,200\}$ as described above. While not all procedures to generate Market Split instances result in feasible problems (e.g., [81], see the study in [5]), we generated them such that feasibility is guaranteed. For $m\leq 7$, we randomly generate problem instances and only include them if a MIP solver can find a feasible solution. For $m>7$, we start with randomly generated feasible solutions and then generate problem instances that have that point as a solution, while resembling the characteristics of the instances in [81]; details are given in Appendix A. In our tests, the resulting instances are still hard for established MIP solvers. Finally, we provide a routine in QOBLIB [193] that, given an instance and a vector $x\in\{0,1\}^{n}$, checks the feasibility of the vector.
@@ 4.1.4 Classical Baseline
We provide modeling examples for an integer linear program and a QUBO, which can be run with any MIP or QUBO solver. For the ILP, we use the constrained formulation with positive slack, minimizing the slack as objective. For the QUBO, we add all constraints as quadratic penalties to the objective. Figure 3 presents a detailed representation of the runtimes needed by Gurobi [95] on a 32-core AMD EPYC-7542 processor using 8 cores and ABS2 [182] on a A100 SXM4 80GB, respectively, for different sizes of instances with different parameters $m$. ABS2 is very fast for smaller instances, but fails to find solutions within one hour for instances larger than size $m=6$. One should note that instances with $m=7$ already appear to be difficult to solve with respect to the required runtime of the solvers. However, in preliminary experiments using enumeration-based methods, either according to [240] or GPU-accelerated [215], we could solve up to size $m=8$ and believe up to size $m=10$ might be feasible with additional engineering effort.
Figure 3: (Market Split) Distribution of solution times using (top) Gurobi with a time limit of 3600s and (bottom) ABS2 with a time limit of 600s for instances with different size $m$ and coefficients range $D$. All trials for $m\geq 8$ timed out for Gurobi and for $m\geq 7$ timed out for ABS2.
@@ 4.2.4 Classical Baseline
Figure 4: (LABS) Solution times using Gurobi and ABS2 for different LABS sequence lengths. Abs. gap denotes (the upper bound on) the absolute gap to the optimal (Gurobi)/best-known (ABS2) solution. Only gaps $>0$ are shown.
@@
Currently, the best-known exact solutions are limited to $N=66$ and $N=127$ for skew-symmetric variants; see [185]. We provide an example where LABS is modeled as a binary quadratically constrained quadratic program (BQCQP), which can be fed to a variety of established solvers. Figure 4 shows results obtained with Gurobi 11.0.0 on an AMD EPYC 7542 32-core processor using 64 threads and a time limit of 2 hours (we report the upper bound on the absolute optimality gap if optimality is not certified), and the results of ABS2 running on a system with two NVIDIA A100-SXM4-80GB GPUs using a time limit of 10 minutes (we report the absolute gap to the best found solution). The largest instance that could be solved to optimality using this approach–although exceeding the usually set time limit–was $N=41$. The respective execution time was 13 hours on a 72-core NVIDIA GraceHopper200 ARM system using Gurobi 11.0.2. An optimal solution was found in 6 hours and 47 minutes, the remaining time was used to prove optimality. Note that problem sizes up to $N=19$ can be solved in less than one second and up to $N=25$ in less than one minute. The optimal computed sequences are shown in Table 3.
@@ 4.3.1 Background
The minimum Birkhoff decomposition problem is a special case of the Birkhoff decomposition problem [63]. In 1946, Birkhoff showed that any $n\times n$ doubly stochastic matrix can be written as a convex combination of permutation matrices [25]. The method of proof used by Birkhoff inspired a class of algorithms for decomposing doubly stochastic matrices known as the Birkhoff algorithm [30, pp. 192], which can decompose a doubly stochastic matrix with at most $(n-1)^{2}+1$ permutation matrices [175]. It is possible to obtain tighter upper bounds on the number of permutations required depending on the number of zeros in the doubly stochastic matrix (see, e.g., [31]); however, the lower bound, i.e., the minimum number of permutations matrices required, is generally not known. In [63], Dufossé and Uçar showed that the problem of finding the minimum Birkhoff decomposition is NP-hard despite the existence of algorithms that can obtain approximate decomposition where the error decreases exponentially with the number of permutations in the decomposition [131].
@@ 4.3.2 Problem statement
Let $D$ be an $n\times n$ doubly stochastic matrix and $P_{i}$ the $i$-th $n\times n$ permutation matrix, $i\in\{1,\dots,n!\}$. Recall a matrix is doubly stochastic if its entries are non-negative and the rows and columns sum to one. Similarly, a permutation matrix is a doubly stochastic matrix with binary entries.
@@
where we define $00^{0}=$ 0, i.e., $|c_{i}|^{0}$ counts the number of non-zero entries, and $s>0$ is a scalar, often equal to one. The goal is to find a subset of permutation matrices such that $D$ is in its convex hull and that this subset contains as few permutation matrices as possible. The following is an example of a decomposition of a $3\times 3$ matrix:
@@ 4.3.3 Instances
The instance set consists of doubly stochastic matrices of size $n\in\{3,\dots,16\}$ generated by sampling $n$ (sparse) and $n^{2}$ (dense) permutations uniformly at random. There are 10 instances for each matrix size and density. The weight associated with each permutation is also sampled uniformly at random with a fixed number of digits after the decimal points. The number of digits after the point depends on the size of the doubly stochastic matrix. We then scaled all coefficients to make them integer. See [193] for a detailed description of how doubly stochastic matrices are generated. By construction, sparse doubly stochastic matrices can be decomposed with at most $n$ permutations, which is a non-trivial upper bound on the minimal decomposition compared to the upper bound $(n-1)^{2}+1$ that applies to all doubly stochastic matrices.
@@ 4.3.4 Classical Baseline
First, we focus on the two convex optimization algorithms devised to obtain sparse approximate decompositions, i.e., not a minimal and exact decomposition. Notably, classical heuristics are discussed for this problem because we also investigate quantum heuristics in Section 5.2. We run the algorithms with their out-of-the-box settings [76, 24] and show the results in Figure 5. The shaded area in Figure 5 (only visible in the sparse case) indicates the maximum and minimum number of permutation matrices required for a given problem size, i.e., the best and worst case instance. The ‘known solution’ corresponds to the number of permutation matrices used to generate the doubly stochastic matrix in the dataset, which is an upper bound to the minimum decomposition. Observe from the figure that Birkhoff+ obtains better performance than Blended FW in both cases. The reason for that is that Birhoff+ is guaranteed to obtain an exact decomposition with at most $(n-1)^{2}+1$ permutations, whereas Blended FW does not have such a property even though it can reduce the approximation error exponentially fast with the number of permutations this adds to the decomposition. In the experiments, we limit the maximum number of permutations of Blended FW to $(n-1)^{2}+1$ and observe decomposition errors in between $10^{-2}$ and $10^{-6}$ depending on the instance. With Birkhoff+, all the decompositions are exact, and their performance varies depending on whether the matrix to decompose is sparse or dense. In the sparse case, Birkhoff+ obtains decompositions that match the number of permutations used to generate the matrices (i.e., $n$) when $n\leq 8$. For $n\geq 9$, the number of permutations required varies significantly depending on the instance (shaded area), especially in the range $n\in[9,12]$. Finally, in the dense case, we obtain that the number of permutations obtained is smaller than the number of permutations used to generate the doubly stochastic matrix (i.e., $n^{2}$). There is no significant variability across instances, and the number of matrices required by Birkhoff+ to obtain an exact decomposition scales linearly up to the problem sizes investigated in this analysis.
@@
Next, we illustrate the performance of CPLEX when solving the following Integer program (IP): $\min_{c_{i},P_{i}}\sum_{i=1}^{k}|c_{i}|^{0}$ s.t. $sD=\sum_{i=1}^{k}c_{i}P_{i}$, $\sum_{i=1}^{k}c_{i}=s$, $c_{i}\in\{0,1,\dots,s\}$, $P_{i}\in\{0,1\}^{n\times n}$, $P_{i}\mathbf{1}=P_{i}^{T}\mathbf{1}=\mathbf{1}$ for all $i\in\{1,\dots,k\}$ where $\mathbf{1}$ is the all-ones vector, and $s$ and $k$ integers. Note that the permutation matrices are formulated as decision variables and all decision variables are represented by integers. Also, note that when the entries of $D$ have a fixed number of decimal points (see Section 4.3.3), we can always select $s$ to be an integer and make $sD$ an integer-valued matrix. For example, if we multiply Eq. (7) by 10, we obtain $\left[\begin{smallmatrix}2&3&5\\ 6&2&2\\ 2&5&3\end{smallmatrix}\right]$, which we can decompose with weights 2, 2, 1, and 5. Parameter $k$ in the formulation controls the decomposition length–which is ideally as small as possible. Thus, to solve the problem, we start the optimization with $k=1$ and increase $k$ by one if CPLEX fails to return a feasible solution. Figure 6 illustrates the runtime of CPLEX as a function of the matrix size. One can see from the figure that the runtime increases rapidly in both cases. With CPLEX on a laptop (M1 Max CPU and 32 GB memory) and a one-hour runtime limit, we can only solve instances up to $n=4$ when the matrix is dense and up to $n=8$ when the matrix is sparse.

### v2

@@ 4 Problem Classes & Instances
Figure 2: The figures plot the problem dimension (MIP: $\sqrt{\#variables\cdot\#constraints}$, QUBO: #variables\#variables) against the density , i.e., the fraction of non-zero coefficients in the objective and, for the MIP instances, constraint functions, of the various problem instances provided in the [202] repository for MIP (top) and QUBO (bottom) formulations of the problem. The legend entries for the various problem classes name the class itself and, in brackets, give the form of the (objective/constraint) function with L and Q denoting linear and quadratic, respectively. Furthermore, the size of the individual points is directly related to the coefficient range of the respective problem instance per problem class–see legend in the center right. The dashed lines illustrate the separation between sparse and dense instances as described in Table 2. Finally, the inset plots illustrate the instances in which the optimal solution is known. It should be noted that while the MIP plot shows all problem instances, the QUBO plot only shows those instances that we could transform into QUBO models with less than $1.6\cdot 10^{5}$ variables, cf. Table 2. The respective model translation is done for Birkhoff, Steiner, Sports, Network, Routing, and Topology with Qiskit Optimization 0.6.1 [201] using the default parameters. Marketsplit, LABS, Portfolio, and Independent Set can be trivially converted to QUBOs using ZIMPL 3.6.1 [141], since there are no slack variables to consider. Note that for Portfolio, this shows the density of the constraints matrix; the quadratic objective is always dense.
@@
The number of decision variables for which MIP or QUBO formulations of the selected instances become difficult ranges from less than 100 to 45,000, and is therefore within reach for testing algorithms on near-term quantum hardware. Depending on the problem, the ranges of the constraint coefficients lie between $1$ and $\approx 10^{7}$, assuming that the coefficients are scaled to integers, as we do throughout the paper. This can pose another challenge for quantum computers as they need to be able to represent the problem data in the necessary resolution.
@@
Figure 2 provides a more detailed overview of the instances provided for the various problem classes, complementing Table 2. For MIP and QUBO formulations of the considered problem instances, it illustrates the number of variables, the problem density, the coefficient range, and whether the optimal solution is known. Please note that if multiple MIP and QUBO models are available (see [202]), we choose to represent the one resulting in the smallest number of variables. We would further like to point out that the QUBO illustration only shows those instances that could be converted from an MIP model into a QUBO model with fewer than $1.6\cdot 10^{5}$ variables. As the plots show, the larger the number of variables, the more likely it is that we do not know the optimal solution. Finally, it can be seen that Market Split and LABS are the two problem classes for which difficult instances are found at the smallest sizes. Independent Set also has hard instances at relatively small sizes; these instances are also sparse and come with a small coefficient range. For these reasons, these three problem classes seem the most amenable for approaching with a quantum algorithm on near-term hardware.
@@
In the remainder of this section, we discuss each problem class, describe the specific instances included, and present baseline results computed with classical optimization solvers. Of course, all instances, used models, and solutions are also made accessible in the [202] repository. We do not claim that our baseline results are the best solutions one can find with a classical computer, but they should provide an indication of what is possible using general solvers without extensive effort. Moreover, the provided MIP and QUBO formulations were not subject to significant optimization efforts but are intended to serve as illustrations. Our goal is to provide a framework to track the progress towards quantum advantage in optimization and have an initial baseline to start with. To build-up expertise and start tracking today’s capabilities, we describe an informed approach to construct warm-up examples from the provided problem instances in Appendix A.
@@ 4.1.1 Background
The market split instances are multi-dimensional subset-sum problems, which were explicitly designed to be hard to solve using classical methods. Since their introduction by [86] in 1999, the progress in solving them has been moderate. Subset-sum is a classic NP-hard problem [90]. The time complexity for enumeration is $\mathcal{O}(2^{n/2}\cdot(n/4))$ [224]. Several other methods have been tried, including smart enumeration, branch-and-cut, and lattice point enumeration with basis reduction [5, 249, 244].
Why is it interesting? Market split problems are a type of multi-dimensional subset-sum problem. This problem class contains various instances for which it is challenging to find feasible solutions with classical digital systems–already for small system sizes, with little more than $100$ variables. Applications of the market split problem could be interesting in various industries, such as the energy sector, e.g., to help derive optimal pricing strategies subject to customer acquisition, retention or competition constraints, and regional regulatory policies [95, 79].
@@ 4.1.2 Problem statement
where a feasible solution exists if the global minimum is zero. With careful choices of the coefficients, this problem becomes hard to solve, e.g., for numbers $m\in\mathbb{N}_{+}$, $D\in\mathbb{N}$, set n=10(m1)n=10(m-1), $I=\{1,\ldots m\},J=\{1,\ldots,n\}$ and $a_{ij}\in\{0,\ldots,D-1\},b_{i}=\left\lfloor\frac{1}{2}\sum_{j\in J}a_{ij}\right\rfloor$ [86].
The best results so far have been achieved using forms of lattice enumeration with basis reduction. [249, 250] report solutions to selected instances up to $m=14$. Traditional branch-and-cut-based MIP solvers often already struggle before size $m\geq 7$. However, they are typically more effective on the lattice reformulation introduced in [5].
@@ 4.1.3 Instances
For this benchmark, we generated four random instances each for $m=\{3,\ldots,15\}$, and $D=\{50,100,200\}$ as described above. While not all procedures to generate Market Split instances result in feasible problems (e.g., [86], see the study in [5]), we generated them such that feasibility is guaranteed. For $m\leq 7$, we randomly generate problem instances and only include them if a MIP solver can find a feasible solution. For $m>7$, we start with randomly generated feasible solutions and then generate problem instances that have that point as a solution, while resembling the characteristics of the instances in [86]; details are given in Appendix C. In our tests, the resulting instances are still hard for classical solvers. Finally, we provide a routine in [202] that, given an instance and a vector $x\in\{0,1\}^{n}$, checks the feasibility of the vector.
@@ 4.1.4 Classical Baseline
We provide modeling examples for an integer linear program and a QUBO, which can be run with any MIP or QUBO solver. For the ILP, we use the constrained formulation with positive slack, minimizing the slack as objective. For the QUBO, we add all constraints as quadratic penalties to the objective. Figure 21 in the Appendix depicts detailed runtimes by Gurobi 11.0.0 [102] and ABS2 [191] for instances with $3\leq m\leq 7$.
Enumeration methods, either Lattice-based according to [249], or Schroeppel-Shamir [224], show better performance and allow solving all instances up to size $m=11$ and partly up to $m=14$. Figure 3 depicts a comparison of the runtimes of the algorithm from [250] on an Intel Xeon E-2288G CPU @ 3.70 GHz and the GPU-accelerated Schroeppel-Shamir algorithm [135] on an Nvidia GH200 system. Note that both methods are capable of computing all solutions to an instance. The time needed to find the “first” solution, however, varies easily by a factor of 10 between the instances.
Figure 3: (Market Split) Comparison of runtimes using a lattice enumeration method and a GPU-accelerated Schroeppel-Shamir algorithm. The colors represent parameter $m$ and the shapes represent parameter $D$.
@@ 4.2.4 Classical Baseline
Figure 4: (LABS) Solution times using Gurobi and ABS2 for different LABS sequence lengths. Abs. gap denotes (the upper bound on) the absolute gap to the optimal (Gurobi)/best-known (ABS2) solution. Only gaps >0>0 are shown.
@@
Currently, the best-known exact solutions are limited to $n=66$ and $n=127$ for skew-symmetric variants; see [194]. We provide an example where LABS is modeled as a binary quadratically constrained quadratic program (BQCQP), which can be fed to a variety of established solvers. Figure 4 shows results obtained with Gurobi 11.0.0 on an AMD EPYC-7542 32-core processor using 64 threads and a time limit of 2 hours (we report the upper bound on the absolute optimality gap if optimality is not certified), and the results of ABS2 running on a system with two Nvidia A100 SXM4 80GB GPUs using a time limit of 10 minutes (we report the absolute gap to the best found solution). The largest instance that could be solved to optimality using this approach–although exceeding the usually set time limit–was $n=41$. The respective execution time was 13 hours on a 72-core Nvidia GH200 system using Gurobi 11.0.2. An optimal solution was found in 6 hours and 47 minutes; the remaining time was used to prove optimality. Note that problem sizes up to $n=19$ can be solved in less than one second and up to $n=25$ in less than one minute. The optimal computed sequences are shown in Table 3.
@@ 4.3.1 Background
The minimum Birkhoff decomposition problem is a special case of the Birkhoff decomposition problem [68]. In 1946, Birkhoff showed that any $n\times n$ doubly stochastic matrix can be written as a convex combination of permutation matrices [28]. The method of proof used by Birkhoff inspired a class of algorithms for decomposing doubly stochastic matrices known as the Birkhoff algorithm [34, pp. 192], which can decompose a doubly stochastic matrix with at most $(n-1)^{2}+1$ permutation matrices [184]. It is possible to obtain tighter upper bounds on the number of permutations required depending on the number of zeros in the doubly stochastic matrix (see, e.g., [35]); however, the lower bound, i.e., the minimum number of permutation matrices required, is generally not known. In [68], Dufossé and Uçar showed that the problem of finding the minimum Birkhoff decomposition is NP-hard despite the existence of algorithms that can obtain approximate decomposition where the error decreases exponentially with the number of permutations in the decomposition [138].
@@ 4.3.2 Problem statement
Let $D$ be an $n\times n$ doubly stochastic matrix and $P_{i}$ the $i$-th $n\times n$ permutation matrix, $i\in\{1,\dots,n!\}$. Recall that a matrix is doubly stochastic if its entries are non-negative and the rows and columns sum to one. Similarly, a permutation matrix is a doubly stochastic matrix with binary entries.
@@
where we define $0^{0}=0$, i.e., $|c_{i}|^{0}$ counts the number of non-zero entries, and $s>0$ is a scalar, often equal to one. The goal is to find a subset of permutation matrices such that $D$ is in its convex hull and that this subset contains as few permutation matrices as possible. The following is an example of a decomposition of a $3\times 3$ matrix:
@@ 4.3.3 Instances
The instance set consists of doubly stochastic matrices of size $n\in\{3,\dots,16\}$ generated by sampling $n$ (sparse) and $n^{2}$ (dense) permutations uniformly at random. There are 10 instances for each matrix size and density. The weight associated with each permutation is also sampled uniformly at random with a fixed number of digits after the decimal points. The number of digits after the point depends on the size of the doubly stochastic matrix. We then scaled all coefficients to make them integers. See [202] for a detailed description of how doubly stochastic matrices are generated. By construction, sparse doubly stochastic matrices can be decomposed with at most $n$ permutations, which is a non-trivial upper bound on the minimal decomposition compared to the upper bound $(n-1)^{2}+1$ that applies to all doubly stochastic matrices. It should be noted that the instances were generated from feasible solutions to guarantee the feasibility of all given problem instances. However, it is not known whether these feasible solutions are optimal.
@@ 4.3.4 Classical Baseline
First, we focus on the two convex optimization algorithms devised to obtain sparse approximate decompositions, i.e., not a minimal and exact decomposition. Notably, classical heuristics are discussed for this problem because we also investigate quantum heuristics in Section 5.2. We run the algorithms with their out-of-the-box settings [81, 27] and show the results in Figure 5. The shaded area in Figure 5 (only visible in the sparse case) indicates the maximum and minimum number of permutation matrices required for a given problem size, i.e., the best and worst case instance. The ‘known solution’ corresponds to the number of permutation matrices used to generate the doubly stochastic matrix in the dataset, which is an upper bound to the minimum decomposition. Observe from the figure that Birkhoff+ obtains better performance than Blended FW in both cases. The reason for that is that Birhoff+ is guaranteed to obtain an exact decomposition with at most $(n-1)^{2}+1$ permutations, whereas Blended FW does not have such a property, even though it can reduce the approximation error exponentially fast with the number of permutations this adds to the decomposition. In the experiments, we limit the maximum number of permutations of Blended FW to $(n-1)^{2}+1$ and observe decomposition errors in between $10^{-2}$ and $10^{-6}$ depending on the instance. With Birkhoff+, all the decompositions are exact, and their performance varies depending on whether the matrix to decompose is sparse or dense. In the sparse case, Birkhoff+ obtains decompositions that match the number of permutations used to generate the matrices (i.e., $n$) when $n\leq 8$. For $n\geq 9$, the number of permutations required varies significantly depending on the instance (shaded area), especially in the range $n\in[9,12]$. Finally, in the dense case, we obtain that the number of permutations obtained is smaller than the number of permutations used to generate the doubly stochastic matrix (i.e., $n^{2}$). There is no significant variability across instances, and the number of matrices required by Birkhoff+ to obtain an exact decomposition scales linearly up to the problem sizes investigated in this analysis.
@@
Next, we illustrate the performance of CPLEX when solving the following Integer program (IP): $\min_{c_{i},P_{i}}\sum_{i=1}^{k}|c_{i}|^{0}$ s.t. $sD=\sum_{i=1}^{k}c_{i}P_{i}$, $\sum_{i=1}^{k}c_{i}=s$, $c_{i}\in\{0,1,\dots,s\}$, $P_{i}\in\{0,1\}^{n\times n}$, $P_{i}\mathbf{1}=P_{i}^{T}\mathbf{1}=\mathbf{1}$ for all $i\in\{1,\dots,k\}$ where $\mathbf{1}$ is the all-ones vector, and $s$ and $k$ integers. Note that the permutation matrices are formulated as decision variables, and all decision variables are represented by integers. Also, note that when the entries of $D$ have a fixed number of decimal points (see Section 4.3.3), we can always select $s$ to be an integer and make $sD$ an integer-valued matrix. For example, if we multiply Eq. (7) by 10, we obtain $\left[\begin{smallmatrix}2&3&5\\ 6&2&2\\ 2&5&3\end{smallmatrix}\right]$, which we can decompose with weights 2, 2, 1, and 5. Parameter $k$ in the formulation controls the decomposition length–which is ideally as small as possible. Thus, to solve the problem, we start the optimization with $k=1$ and increase $k$ by one if CPLEX fails to return a feasible solution. Figure 6 illustrates the runtime of CPLEX as a function of the matrix size. One can see from the figure that the runtime increases rapidly in both cases. With CPLEX on a laptop (M1 Max CPU and 32 GB memory) and a one-hour runtime limit, we can only solve instances up to $n=4$ when the matrix is dense and up to $n=8$ when the matrix is sparse.